OpenStudy (anonymous):

Randy wants to move his savings account to a new bank that pays a better interest rate of 3.5% so that he can earn $100 in interest faster than at his old bank. If he moves $800 to the new bank, how long will it take for him to earn the $100 in interest?

OpenStudy (amistre64):

depends on how the interest is calculated, compounded so many times a year

OpenStudy (amistre64):

assuming simple interest is not a real life application, but may be plausible for the question

OpenStudy (amistre64):

800(1+r/n)^(nt) = 900 if compounded n times a period

OpenStudy (anonymous):

so I just do that?

OpenStudy (amistre64):

i would ..... bu tthe question is ambiguous since we dont know what sort of interest or compounding factor we have to work with

OpenStudy (amistre64):

we can clove for t, but it will be in terms of n; or if simple interest 100 = 800rt, and solving for t is pretty easy but we know nothing about the type of interest to determine which method they desire

OpenStudy (amistre64):

**we can solve for t ....

OpenStudy (anonymous):

Oh ok

OpenStudy (amistre64):

P(1+r/n)^(nt) = A (1+r/n)^(nt) = A/P nt log(1+r/n) = log(A/P) t = log(A/P) / [n log(1+r/n)] so if n=1, 4, 12, etc .. we can determine a time value for it using compound interest

OpenStudy (anonymous):

I found it it's supposed to be simple

OpenStudy (amistre64):

simple is much simpler then :) 100 = 800(.035)t divide off the excess

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you